### v1

# Learning to Learn Transferable Generative Attack for Person Re-Identification

Abstract
Deep learning-based person re-identification (re-id) models are widely employed in surveillance systems and inevitably inherit the vulnerability of deep networks to adversarial attacks. Existing attacks merely consider cross-dataset and cross-model transferability, ignoring the cross-test capability to perturb models trained in different domains. To powerfully examine the robustness of real-world re-id models, the Meta Transferable Generative Attack (MTGA) method is proposed, which adopts meta-learning optimization to promote the generative attacker producing highly transferable adversarial examples by learning comprehensively simulated transfer-based cross-model&dataset&test black-box meta attack tasks. Specifically, cross-model&dataset black-box attack tasks are first mimicked by selecting different re-id models and datasets for meta-train and meta-test attack processes. As different models may focus on different feature regions, the Perturbation Random Erasing module is further devised to prevent the attacker from learning to only corrupt model-specific features. To boost the attacker learning to possess cross-test transferability, the Normalization Mix strategy is introduced to imitate diverse feature embedding spaces by mixing multi-domain statistics of target models. Extensive experiments show the superiority of MTGA, especially in cross-model&dataset and cross-model&dataset&test attacks, our MTGA outperforms the SOTA methods by 21.5% and 11.3% on mean mAP drop rate, respectively. The code of MTGA will be released after the paper is accepted.
Index Terms:
Re-id, Transferable Adversarial Example, Meta-learningI Introduction
Person re-identification aims at retrieving specific persons from security surveillance systems [1, 2]. Along with the advancement of deep neural networks[3, 4, 5], it has made remarkable progresses and been widely applied to intelligent surveillance systems [6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. However, it has been found that deep neural networks are vulnerable to adversarial attacks[16, 17, 18, 19, 20], which can mislead deep neural network models by adding imperceptible perturbations to benign images. Deep learning-based re-id models inevitably inherit the vulnerability of deep networks to adversarial samples[21, 22], which makes public safety under great threat. To study the security of surveillance systems, it is important to explore the vulnerability of the deep learning-based re-id models to adversarial samples.
Recently, some works [23, 21, 24, 22] have demonstrated that re-id models are susceptible to adversarial examples and introduced white-box adversarial metric attack methods to attack re-id models. These methods are not suitable in realistic scenarios, where parameters of target re-id models are not accessible. Transferable adversarial examples against black-box re-id models are then studied [25, 26, 27, 28, 29]. Different from transfer-based black-box attacks for classification tasks, which assume attackers have access to the training data of target model and generally only consider cross-model transferability among models trained in the same data distribution[30, 31], attacks on black-box re-id models are more challenging due to the cross-model (architecture discrepancy between surrogate model and target model), cross-dataset (domain discrepancy between training image and target image) and cross-test (domain discrepancy between target image and target model) transfer capabilities are supposed, like Fig. 1 shows. Specifically, re-id is an open-set task [32, 33], where identities in the training and testing sets are non-overlapped and unseen query images often encounter a large domain shift [34], thus cross-dataset transferability is necessary for black-box adversarial attacks against re-id models. Except for cross-model transferability to attack models with different architectures, cross-test capability should take into account to attack models with different feature embedding spaces, since target re-id models could be trained with arbitrary domain datasets. However, existing transfer-based re-id attacks do not fully consider these aspects, either ignoring cross-dataset capabilities[27, 28] or merely focusing on cross-model transferability and neglecting the cross-test capabilities[26, 29, 25], which leads to insufficient transferability of generated adversarial samples to effectively test the robustness of real-world re-id models.
In order to generate highly transferable adversarial examples against person re-id models, we propose the Meta Transferable Generative Attack (MTGA) approach, which utilizes meta-learning optimization to guide the generative attacker possessing the generic transferability by learning multiple simulated cross-model&dataset&test black-box meta attack tasks. Various train-test processes of cross-model&dataset transfer-based black-box attacks are first generated as meta-learning tasks by Cross-model&dataset Attack Simulation (CAS) method. In terms of cross-dataset mimicking, multi-source datasets in the data zoo are utilized to randomly represent the adversarial attack training data and unseen domain testing data. For cross-model imitation, the agent model and the target model are picked differently in model zoo, which consists of three classical re-id models that can well represent global-based, part-based and attention-based approaches, considering these three types of re-id methods are most widely applied. Besides, considering limited surrogate model resources for constructing meta-attack tasks and given the observation that different models focus on different discriminative regions in recognition[35], the Perturbation Random Erasing (PRE) module is introduced to erase randomly selected perturbation regions to prevent the attacker from only learning to destroy the model-specific features or salient features, thus enhance the cross-model generalization of adversarial examples. Meanwhile, the Normalization Mix (NorMix) strategy is devised to mimic cross-test embedding spaces by dynamically mixing the multi-domain batch-norm statistics of the target model, boosting attackers learning the ability of attacking target models that trained in different domain data. Extensive experiments on numerous re-id benchmarks and models show our MTGA achieves state-of-the-art (SOTA) transferability on all six black-box attack scenarios, demonstrating the effectiveness of our method. Especially for cross-model&dataset and cross-model&dataset&test attack, our MTGA surpasses the SOTA methods by 21.5% and 11.3% on mean mAP drop rate, respectively. In summary, our main contributions are as follows:
We propose a novel Meta Transferable Generative Attack (MTGA) method that creates extensive cross-model&dataset&test black-box meta attack tasks for adversarial generative attackers to learn to generate more generic and transferable adversarial examples against real-world re-id models.
Cross-model&dataset Attack Simulation approach is presented to mimic transfer-based cross-model and cross-dataset meta attack tasks by selecting distinct model and dataset for meta-train and meta-test processes.
Perturbation Random Erasing module is devised to enhance the transferability by suppressing the model-specific features corruption and encouraging disruption of entire feature rather than only discriminative feature.
Normalization Mix strategy is introduced to simulate cross-test attack by dynamically mixing the multi-domain batch-norm statistics of the target model, diversifying feature embedding spaces of re-id models.
II Related Works
II-A Transferable Adversarial Attack
Szegedy et al.[16] found the intriguing transferability of attack examples, which permits attackers to generate adversarial example from surrogate models to attack black-box target models. Since this property of adversarial examples poses real-world DNN applications under serious security concerns, there have been extensive works aiming to improve the transferability of adversarial examples, which can be categorized into four groups, namely input transformation attacks [35, 36, 37, 38], gradient modification attacks [39, 40, 41], intermediate feature attacks [42, 43, 44] and model ensemble attacks [45, 46, 47]. Nevertheless, these approaches merely consider the cross-model transferability, presuming that the distribution of the attacked images and the target model training data are consistent, which can not been guaranteed in real-world situations. There are very few literatures focusing on this issue. Naseer et al. [48] trained a generative network that produces transferable cross-dataset perturbations by maximizing the fooling gap using relativistic supervisory signal. Zhang et al. [31] trained the attacker to disrupt low-level features and enhanced the transferability towards black-box domains by randomly normalizing benign images at image-level. Li et al. [30] trained a domain-agnostic feature extractor by self-supervised learning as the surrogate model to accomplish the cross-dataset transferability. Our MTGA is proposed for more complicated cross-model&dataset&test transferable attacks on black-box re-id models.
II-B Adversarial Attack Against Person Re-id
Previous attack methods [49, 50, 51] have mainly concentrated on the image classification task, aiming to significantly alter class predictions. But these methods are inapplicable to attack re-id models[23], since the re-id task is an open-set task. To effectively attack re-id models, Bai et al. [21], Zheng et al. [23] and Bouniot et al. [24] proposed different metric attack methods based on feature similarity calculation. Nevertheless, these works conducted white-box attacks, which need to know the parameters of target re-id models and are only valid for seen adversarial attack training data. They are not practical, because attackers need to attack unknown models and unseen queries in the realistic scenarios. To accomplish black-box attacks against re-id models, some researches [25, 27, 26, 28, 29] studied the transferable attacks for re-id system. Yang et al. [26] and Subramanyam [29] only enhanced the cross-dataset transferability by adopting multi-source datasets in additive and generative attack, respectively. Wang et al. [28] also ignored the cross-model transfer capabilities and developed a multi-stage discriminator network for cross-dataset general attack learning. Ding et al. [27] merely focused on cross-model transferability, introducing a model-insensitive regularization term for universal attack against different CNN structures. Yang et al. [25] built a combinatorial attack that consists of a functional color attack and universal additive attack to promote the cross-model&dataset of the attack. However, they attacked domain-specific target models, neglecting the diversity of feature embedding spaces of the same model. These existing transfer-based re-id black-box attacks have not yet achieved high transferability, without take into account the cross-test transferability of re-id adversarial examples.
II-C Meta-learning
Meta-learning is a learning-to-learn[52] algorithm, which aims to improve further learning performance by distilling the experience from multiple learning episodes (i.e. meta-train and meta-test processes)[53, 54]. It has been widely used in deep learning tasks, e.g. few-shot learning, domain generalization and hyperparameter optimization. Recently, some meta-learning based transferable adversarial attack methods have been proposed, the underlying concept of which is to construct numerous meta transfer attack tasks. Yuan et al. [55] enhanced the cross-model transferability by composing different cross-model meta attack tasks. Fang et al. [56] composed transfer attack tasks with data augmentation and model augmentation, through randomized data transformation and model backpropagation altering. Yin et al. [57] generalized the generic prior of examples by treating attack on each examples as one task and fine-tuning the surrogate model during the meta-test process.
Distinct from above adversarial attack methods for re-id and meta-learning based attack methods, our method constructs extensive cross-model&dataset&test black-box adversarial attack tasks for attackers to learn how to generate more generic and transferable adversarial examples. And our CAS, PRE and NorMix modules are quite distinct from others.
III Methodology
In this section, we first present the problem definition of the generative adversarial attack against re-id models in Section III-A. The overall framework of MTGA and the meta-learning optimization is then introduced in Section III-B. Right after that, the details about how to generate extensive transfer-based black-box meta attack tasks are described in Section III-C. Finally, the optimization procedure of our method are given in Section III-D.
III-A Problem Definition
The goal of our proposed MTGA is to optimize the parameters of the adversarial generator to produce adversarial perturbation for each benign image . The adversarial example is produced by adding additive perturbation to the query image to attack the re-id models for outputting incorrect retrieval images. To ensure adversarial perturbations are imperceptible, the maximum magnitude of perturbations allowed to be added cannot exceed .
The adversarial generator is first trained in the white-box way, knowing the attacked queries and the target re-id model. Then, it is fixed and used to produce perturbations for unseen data to attack black-box re-id models.
III-B Overall Framework
The proposed MTGA is based on the meta-learning optimization framework, as Fig. 2 shows. Meta tasks are generated to simulate the train-test processes of transfer-based black-box attack to train the generative attacker learning to produce generic adversarial examples. The data zoo and model zoo that contain multiple datasets and multiple re-id models are first prepared for meta-task generation. In each meta task , datasets and re-id models for meta-train () and meta-test () processes are distinctly selected from the data zoo and model zoo to mimic training data and unseen test data, as well as the surrogate model and target model. The discriminator is adopted in optimization processes to distinguish the adversarial images from benign images to boost generator producing deceptive perturbations. The parameters of generator are updated after meta-train process. Then, in the meta-test process, generates adversarial perturbations for with the updated to test the transferability of trained generator. The perturbations are randomly erased by the PRE strategy and the features are projected to diverse embedding spaces through the NorMix module by mixing the and feature distributions that extracted by . The meta-test errors of generated tasks serves as the training error of the various transfer-based black-box attack processes to optimize the adversarial generator.
III-C Meta Task Generation
The meta-task consists of a meta-training and a meta-testing process. Meta-train process plays the role of transfer-based black-box attack training process, which utilizes white-box agent models and selected data to train the adversarial generator. And the meta-test process plays the role of transfer-based black-box attack testing process, which tests the transferability of the trained attacker against black-box target model and unseen images. By learning from generated black-box attack tasks, attackers can learn how to generate adversarial examples to attack black-box re-id models. In terms of better learning for generating transferable and generalizable perturbations, a large number of meta-tasks that take all variations of realistic transfer-based black-box attacks into account should be constructed. Specifically, our approach generates diverse cross-model&dataset&test attack tasks by performing the following three methods.
Cross-model&dataset Attack Simulation method. Because of the unknown parameters of the re-id model and unseen domain queries to be attacked in black-box scenarios, the adversarial generator needs to learn to handle the cross-model and cross-dataset attack situations. To mimic this case, Cross-model&dataset Attack Simulation method is proposed, which makes the target model and input data different during meta-train and meta-test process. Concretely, the data zoo and the model zoo that contains multiple datasets and multiple re-id models are constructed, from which CAS randomly selects distinct models and data for meta-train and meta-test processes to simulate cross-model and cross-dataset attacks. To represent numerous models well, CAS takes baseline models of three mainstream approaches (i.e. global-based, part-based and attention-based) to construct the model zoo.
Perturbation Random Erasing strategy. Although there are several surrogate models in the model zoo to allow the attacker learning to handle cross-model attack scenarios, the number of these models is still limited, which may result in the attacker only learning to attack model-specific features. To address this problem, the Perturbation Random Erasing strategy is proposed to boost the attacker to disrupt holistic person features. PRE randomly selects a rectangle region of generated perturbations and erases it to generate incomplete perturbations. These incomplete perturbations prompt the attacker not to rely only on corrupting specific region features, as perturbations in these specific regions may be erased, leading to the failure of damaging specific region features. PRE is adopted in the meta-test process to test the attack error of trained adversarial attackers with generated incomplete perturbations, optimizing that error will enhance the attacker to achieve holistic destruction of image features and improve the transferability against black-box models.
Normalization Mix module. The models that trained with different domain data could project person images to various feature embeddings, even though they share the same model architecture. NorMix is devised to project features to different feature embedding spaces, which is applied in meta-test process to promote the attacker learning to handle this cross-test issue. Specifically, there are multiple batch-norm layers [58] across the re-id model architectures, the statistics of which imply the distribution of the model training data. The batch normalization is formulated as
where is the input feature, and are the mean and variance of , and are learnable affine parameters used for linear transformation. To get diverse feature embeddings that the test data may be projected by the target model, the statistic of each batch-norm layer is mixed by
where and are the empirical mean and variance of the pretrained meta-test model , and are the training statistics of the and is the mix coefficient that sampled from Beta Distribution. With mixed mean and variance , meta-test data features can be embedded to different feature spaces by
where and are copied from the batch-norm layers of meta-test model.
III-D Optimization Procedure
The parameters of adversarial generator are supposed to be optimized by the meta-learning optimization. To disrupt the retrieval list of generated adversarial examples, the attacked image features should be far away from the original features. In our MTGA, the adversarial Euclidean Distance () loss
is applied to corrupt the similarity of adversarial features and benign features. Meanwhile, and are trained by the GAN loss respectively, denote as:
Meta-train. With the and , the objective function of meta-train process is calculated by
Meta-test. After meta-train process, the parameters of is updated to , and meta-test loss is expressed by
Meta Optimization. The final loss consists of the meta-test errors for each meta-task, formulated as
which represents the error of adversarial generator with parameters for different cases of transfer-based black-box attacks. By optimizing the , adversarial generator that produces highly transferable adversarial examples against different black-box re-id models can be learned. The optimization procedure is summarized in Algorithm 1.
IV Experiments
IV-A Experimental Setup
Evaluation settings. To verify the attack performance of our methods against real-world re-id models, we comprehensively consider different adversarial attack scenarios and set up six attack settings. The details of these settings are showed in Tab. I. The cross-model attack setting implies the black-box target model architecture is different with the surrogate model, yet the training domain of them is the same. The cross-dataset attack setting means the domain of query images and re-id models are different from the white-box attack training process, and query images and the target model training data are in the same domain. These settings are the same as transfer-based black-box re-id attacks proposed by [25], to which we have added cross-test setting. The cross-test setting indicates that the domains of the query data and the target model are different, simulating the most practical application of the real-world re-id models.
Model zoo and data zoo. Model zoo is composed of IDE[2], PCB[59] and ViT[13], which are all trained on the DukeMTMC[60] datasets. And the data zoo consists of DukeMTMC[60], CUHK03[61], and MSMT17[62] datasets.
Black-box re-id models and unseen queries. To evaluate the transferability of our adversarial generator to different re-id models, numerous re-id models (i.e. BOT[63], LSRO[64], MuDeep[65], Aligned[66], MGN[67], HACNN[68], Transreid[13], PAT[69]) are taken to act as the black-box re-id models. Notably, these models are in different backbones, including ResNet[3] (i.e. BOT[63]), ViT[4] (i.e. Transreid[13], PAT[69]), DenseNet[70] (i.e. LSRO[64]) and Inception-v3[71] (i.e. MuDeep[65]). Also, these models are in different architectures, including global-based (i.e. BOT[63]), part-based (i.e. MGN[67]) and attention-based (i.e. HACNN[68]). In order to test the transferabilities on different domain models, these models are trained on different domain datasets (i.e. Market[72] and DukeMTMC[60]). Meanwhile, to test the transferability of our attacker to unseen queries, VIPeR[73] and Market[72] datasets play the role of unseen domain data.
| Attack Settings |
| Cross-dataset | ✘ | ✔ | ✘ | ✔ | (Duke) | Market | (Market) | |||||||||||||||||
| Cross-dataset&test | ✘ | ✔ | ✘ | ✘ | (Duke) | VIPeR | (Market) | |||||||||||||||||
| Cross-model | ✔ | ✘ | ✔ | ✔ | (Duke) | Duke | (Duke) | |||||||||||||||||
| Cross-model&test | ✔ | ✘ | ✘ | ✘ | (Duke) | Duke | (Market) | |||||||||||||||||
| Cross-model&dataset | ✘ | ✘ | ✘ | ✔ | (Duke) | Market | (Market) | |||||||||||||||||
| Cross-model&dataset&test | ✘ | ✘ | ✘ | ✘ | (Duke) | VIPeR | (Market) |
| Methods | IDE | PCB | ViT | aAP | mDR |
| None | 75.5 | 70.7 | 86.5 | 77.6 | - |
| MetaAttack | 4.2 | - | - | - | - |
| Mis-Ranking | 26.9 | - | - | - | - |
| MUAP | 19.3 | - | - | - | - |
| MetaAttack* | 20.2 | 35.8 | 61.1 | 39.0 | 49.7 |
| Mis-Ranking* | 16.8 | 36.8 | 48.4 | 34.0 | 56.1 |
| MUAP* | 14.0 | 26.0 | 42.1 | 27.4 | 64.7 |
| MTGA* | 17.1 | 26.6 | 43.7 | 29.1 | 62.5 |
| MTGA(Ours) | 10.8 | 25.5 | 38.4 | 24.9 | 67.9 |
| Methods | IDE | PCB | ViT | aAP | mDR |
| None | 30.0 | 33.0 | 51.0 | 38.0 | - |
| MetaAttack | 10.0 | - | - | - | - |
| Mis-Ranking | 14.2 | - | - | - | - |
| MUAP | 11.9 | - | - | - | - |
| MetaAttack* | 14.1 | 24.7 | 40.7 | 26.5 | 30.3 |
| Mis-Ranking* | 12.4 | 25.9 | 34.4 | 24.2 | 36.2 |
| MUAP* | 11.9 | 20.4 | 35.9 | 22.7 | 40.2 |
| MTGA* | 12.7 | 22.4 | 33.0 | 22.7 | 40.3 |
| MTGA(Ours) | 10.4 | 21.9 | 30.7 | 21.0 | 44.7 |
| Methods | Global-based | Part-based | Attention-based | aAP | mDR | |||||
| BOT | LSRO | MuDeep | Aligned | MGN | HACNN | Transreid | PAT | |||
| None | 76.2 | 55.0 | 43.0 | 69.7 | 66.2 | 60.2 | 79.6 | 70.6 | 65.0 | - |
| MetaAttack | 14.9 | 44.0 | 31.8 | 49.5 | 57.4 | 54.6 | 75.3 | 64.5 | 49.0 | 24.6 |
| Mis-Ranking | 14.4 | 6.8 | 8.0 | 16.5 | 8.4 | 8.8 | 34.5 | 42.9 | 17.5 | 73.1 |
| MUAP | 16.3 | 9.2 | 11.1 | 23.1 | 11.4 | 13.8 | 34.2 | 40.4 | 19.9 | 69.4 |
| MetaAttack* | 23.2 | 15.0 | 11.7 | 22.9 | 13.6 | 19.6 | 43.6 | 40.8 | 23.8 | 63.4 |
| Mis-Ranking* | 6.8 | 2.0 | 9.9 | 8.7 | 4.3 | 6.6 | 16.3 | 22.3 | 9.6 | 85.2 |
| MUAP* | 18.6 | 8.2 | 8.5 | 16.5 | 7.0 | 11.4 | 29.9 | 32.0 | 16.5 | 74.6 |
| MTGA* | 7.9 | 3.1 | 7.8 | 8.7 | 4.4 | 4.9 | 15.0 | 23.2 | 9.4 | 85.5 |
| MTGA(Ours) | 5.1 | 1.4 | 7.2 | 6.5 | 3.2 | 4.9 | 13.8 | 19.9 | 7.7 | 88.2 |
| Methods | Global-based | Part-based | Attention-based | aAP | mDR | |||||
| BOT | LSRO | MuDeep | Aligned | MGN | HACNN | Transreid | PAT | |||
| None | 14.9 | 13.5 | 4.5 | 18.3 | 22.3 | 11.2 | 43.6 | 44.6 | 21.6 | - |
| MetaAttack | 4.9 | 11.8 | 4.3 | 12.6 | 19.9 | 10.8 | 41.3 | 40.1 | 18.2 | 15.7 |
| Mis-Ranking | 9.2 | 6.4 | 2.1 | 9.9 | 11.3 | 5.5 | 29.3 | 35.4 | 13.6 | 37.0 |
| MUAP | 7.2 | 5.9 | 2.6 | 10.4 | 10.4 | 6.0 | 28.4 | 31.9 | 12.9 | 40.3 |
| MetaAttack* | 6.5 | 5.5 | 2.9 | 8.7 | 10.1 | 6.4 | 30.1 | 31.2 | 12.6 | 41.7 |
| Mis-Ranking* | 6.7 | 4.5 | 2.3 | 8.3 | 7.9 | 4.0 | 22.0 | 26.5 | 10.3 | 52.3 |
| MUAP* | 5.0 | 3.5 | 2.3 | 8.5 | 7.5 | 4.8 | 22.5 | 24.7 | 9.9 | 54.4 |
| MTGA* | 6.7 | 4.8 | 1.9 | 7.5 | 7.6 | 3.4 | 20.6 | 25.4 | 9.7 | 55.1 |
| MTGA(Ours) | 5.5 | 3.4 | 1.9 | 7.0 | 6.3 | 3.4 | 18.7 | 23.6 | 8.7 | 59.7 |
Evaluation metrics. The adversarial attack performance of the generated adversarial samples against different re-id models is measured by three metrics, mean Average Precision (mAP)[72], average mAP (aAP) and mean mAP Drop Rate (mDR)[27]. The aAP is calculated by
where represents mAP of the -th re-id models. The mDR is designed to show the success rate of the adversarial attacks to multiple re-id models and is formulated as
where is the aAP of the re-id models on the benign images and is on the generated adversarial examples.
Implementation Details. MAML[54] is adopted as our meta-learning framework and in each iteration meta-tasks are generated. Adam[74] optimizer is employed to optimize the model parameters. The learning rate of inner loop and outer loop are set to 1e-4 and 2e-4. The generator and discriminator model are referenced to the Mis-Ranking[28]. All experiments are performed by -bounded attacks with , where is the upper bound for the change of each pixel. The mix coefficient of NorMix is sampled from Beta Distribution, i.e., .
IV-B Comparison with State-of-the-art Methods
We compare our proposed MTGA method with state-of-the-art attack methods on transferable black-box re-id attacks, including MUAP[27], Mis-Ranking[28], MetaAttack[25]. These methods are all re-trained by attacking IDE[2] on DukeMTMC[60]. Unlike other methods, MetaAttack[25] method incorporates the color attack in addition to the additive perturbation. For a fair comparison, we only compare the attack performances of its additive perturbation. Meanwhile, based on these original methods, we train MetaAttack*, Mis-Ranking*, MUAP* and MTGA* in the ensemble training setting by attacking models in the model zoo (i.e. IDE[2], PCB[59] and ViT[13]) with dataset in data zoo (i.e. DukeMTMC[60], CUHK03[61], and MSMT17[62]). The experiment details of training data, surrogate model, target data and target model are shown in Tab. I. The comparison results on the mAP, aAP and mDR of six black-box attack settings are shown in Tab. III to Tab. VII.
Comparisons with original SOTA methods. It can be seen that in every black-box attack scenario, our MTGA performs much better than other SOTA methods on attacking multiple black-box re-id models. For most practical and challenging cross-model&dataset&test scenario, our MTGA achieves a superior performance of 18.5% aAP and 51.3% mDR score, which outperforms the SOTA methods by 4.3% and 11.3% in terms of aAP and mDR. For cross-model&dataset attack setting, our MTGA also gets the best transferability results, surpassing others by 16.5% and 21.5% in terms of aAP and mDR.
| Methods | Global-based | Part-based | Attention-based | aAP | mDR | |||||
| BOT | LSRO | MuDeep | Aligned | MGN | HACNN | Transreid | PAT | |||
| None | 85.4 | 77.2 | 49.9 | 79.1 | 82.1 | 75.2 | 86.6 | 78.4 | 76.7 | - |
| MetaAttack | 26.3 | 68.6 | 37.8 | 59.4 | 73.0 | 63.9 | 80.0 | 67.7 | 59.6 | 22.3 |
| Mis-Ranking | 46.3 | 36.7 | 11.9 | 47.5 | 46.7 | 27.0 | 65.2 | 63.4 | 43.1 | 43.8 |
| MUAP | 42.9 | 35.7 | 9.7 | 48.0 | 40.6 | 23.8 | 58.3 | 59.7 | 39.8 | 48.1 |
| MetaAttack* | 38.5 | 36.5 | 18.3 | 38.0 | 44.0 | 32.6 | 62.7 | 55.0 | 40.7 | 46.9 |
| Mis-Ranking* | 33.9 | 23.0 | 11.2 | 36.5 | 32.3 | 18.1 | 47.6 | 48.6 | 31.4 | 59.1 |
| MUAP* | 28.7 | 19.5 | 10.3 | 36.0 | 28.5 | 20.4 | 44.0 | 45.6 | 29.1 | 62.0 |
| MTGA* | 31.1 | 21.8 | 8.8 | 31.3 | 27.8 | 13.8 | 42.6 | 43.6 | 27.6 | 64.0 |
| MTGA(Ours) | 24.3 | 14.2 | 6.2 | 27.7 | 24.0 | 11.5 | 37.9 | 40.5 | 23.3 | 69.6 |
| Methods | Global-based | Part-based | Attention-based | aAP | mDR | |||||
| BOT | LSRO | MuDeep | Aligned | MGN | HACNN | Transreid | PAT | |||
| None | 32.7 | 33.5 | 25.8 | 35.3 | 35.8 | 29.0 | 56.2 | 56.0 | 38.0 | - |
| MetaAttack | 16.4 | 30.0 | 22.5 | 28.2 | 34.1 | 26.1 | 53.6 | 50.9 | 32.7 | 13.9 |
| Mis-Ranking | 19.1 | 16.7 | 12.1 | 20.5 | 24.3 | 15.8 | 41.1 | 46.6 | 24.5 | 35.5 |
| MUAP | 18.3 | 14.1 | 12.4 | 22.6 | 20.1 | 15.5 | 36.1 | 43.4 | 22.8 | 40.0 |
| MetaAttack* | 19.1 | 21.3 | 17.5 | 23.1 | 24.9 | 19.3 | 45.4 | 45.0 | 27.0 | 28.9 |
| Mis-Ranking* | 18.2 | 13.4 | 13.8 | 20.4 | 18.4 | 13.6 | 34.7 | 38.6 | 21.4 | 43.7 |
| MUAP* | 18.3 | 15.2 | 13.6 | 24.6 | 21.4 | 16.1 | 38.5 | 40.8 | 23.6 | 38.0 |
| MTGA* | 16.1 | 13.0 | 11.9 | 20.6 | 18.5 | 11.6 | 31.2 | 39.0 | 20.2 | 46.8 |
| MTGA(Ours) | 14.9 | 10.3 | 9.6 | 18.9 | 15.8 | 10.8 | 31.3 | 36.1 | 18.5 | 51.3 |
| Methods | Global-based | Part-based | Attention-based | aAP | mDR | |||||
| BOT | LSRO | MuDeep | Aligned | MGN | HACNN | Transreid | PAT | |||
| None | 85.4 | 77.2 | 49.9 | 79.1 | 82.1 | 75.2 | 86.6 | 78.4 | 76.7 | - |
| Baseline | 46.9 | 38.3 | 18.5 | 53.8 | 51.0 | 26.7 | 68.4 | 63.1 | 45.8 | 40.2 |
| +CAS | 30.9 | 19.4 | 7.3 | 29.1 | 28.9 | 13.5 | 44.4 | 44.5 | 27.2 | 64.5 |
| +PRE | 27.5 | 16.3 | 7.7 | 29.1 | 25.6 | 13.8 | 41.8 | 42.8 | 25.5 | 66.7 |
| +NorMix | 24.3 | 14.2 | 6.2 | 27.7 | 24.0 | 11.5 | 37.9 | 40.5 | 23.3 | 69.6 |
Comparisons with ensemble trained SOTA methods. Although the transferability of the ensemble trained SOTA methods is better than the corresponding original methods, our MTGA still performs better than the SOTA methods that use the resources of our model zoo and data zoo for ensemble training. The superiority of our MTGA than ensemble training methods can be observed in Tab. III to Tab. VII. Specifically, for complicated cross-model&dataset and cross-model&dataset&test black-box attack, our MTGA surpasses them by 7.6% and 7.6% on mDR, respectively.
IV-C Ablation Studies
The ablation study results of CAS, PRE and NorMix modules are presented in Tab. VIII. The baseline model is trained without meta-learning scheme. It uses IDE(DukeMTMC) as the surrogate model and utilizes the DukeMTMC[60] benchmark as training data to train the adversarial generator. Ablation experiments are tested on cross-model&dataset black-box attack case.
The effectiveness of CAS. It can be observed that the incorporation of CAS module results in a significant decrease of 18.6% in aAP and an increase of 23.7% in mDR, which proves the effectiveness of proposed CAS module. The considerable increase in the transferability of the generated adversarial examples illustrates that the CAS module is able to simulate the black-box transfer-based attack tasks very well.
The effectiveness of PRE. Tab. VIII shows the advantage of PRE module, where aAP decreases from 27.2% to 25.5% and mDR increases from 64.5% to 66.7% after the PRE module is added into the training. Also, the Grad-CAM[75] visualization in Fig. 3 shows that PRE can effectively prevent the attacker from learning to corrupt model-specific features. Concretely, Fig. 3a shows that models with different architectures concentrate on different part of persons. And Fig. 3b reflects that without the PRE module, generated adversarial examples merely mislead models to concentrate on different person part features, which results in poor transferability of attacks. Moreover, the attention maps in Fig. 4c demonstrate that the PRE module promotes the holistic feature corruption of person images, enhancing the transferabilities of adversarial examples.
The effectiveness of NorMix. The NorMix module maps the data to diverse feature subspaces, promoting the attacker to be effective not only in the feature subspace of the training models. It is seen in Tab. VIII that the NorMix module improves the mDR from 66.7% to 69.6%, which shows the effectiveness of our NorMix module.
The effectiveness of discriminator. The discriminator is a kind of defence model that recognizes AEs generated from various domains and models, whose feedback helps attackers to generate more transferable AEs. Tab. IX shows a degradation of attack performance without discriminator, demonstrating its effectiveness.
| Methods | Global-based | Part-based | Attention-based | aAP | |||||
| BOT | LSRO | MuDeep | Aligned | MGN | HACNN | Transreid | PAT | ||
| None | 85.4 | 77.2 | 49.9 | 79.1 | 82.1 | 75.2 | 86.6 | 78.4 | 76.7 |
| w/ D | 24.3 | 14.2 | 6.2 | 27.7 | 24.0 | 11.5 | 37.9 | 40.5 | 23.3 |
| w/o D | 25.0 | 15.9 | 7.4 | 29.9 | 26.2 | 13.4 | 40.5 | 43.3 | 25.2 |
| Methods | Global-based | Part-based | Attention-based | aAcc | |||||
| BOT | LSRO | MuDeep | Aligned | MGN | HACNN | Transreid | PAT | ||
| CWA | 43.1 | 57.1 | 92.2 | 39.3 | 54.5 | 57.1 | 43.2 | 47.9 | 54.3 |
| AdaEA | 47.4 | 54.1 | 88.8 | 42.4 | 52.6 | 54.2 | 49.7 | 46.3 | 54.4 |
| NTKL | 68.6 | 37.2 | 76.1 | 55.9 | 44.9 | 40.4 | 52.4 | 55.2 | 53.8 |
| Ours | 33.4 | 9.8 | 52.0 | 23.2 | 15.4 | 9.9 | 33.5 | 32.3 | 26.2 |
The effectiveness of meta-learning. The comparisons between MTGA (trained in meta-learning way) and MTGA* (trained in ensemble-learning way) in the Tab. III to Tab. VII show that MTGA performs much better than MTGA*, which demonstrates the effectiveness of the meta-learning optimization in our method. For example, in cross-dataset and cross-dataset&test settings, MTGA outperforms MTGA* by 5.4% and 4.4% mDR, respectively. The advantage of meta-learning optimization is that it learns to possess transferability capabilities by learning meta tasks, rather than get the optimal solution to the learning resources.
| Methods | MetaAttack | MUAP | Mis-Rank | Ours |
| SSIM | 0.838 | 0.948 | 0.951 | 0.935 |
To further verify the effects of meta-learning and eliminate the effects of data zoo and model zoo, we compare with SOTA classification ensemble attacks (i.e. CWA [76], AdaEA [77], NTKL [78]). Since they only integrate multiple models without using multiple datasets, we retrained a model without the data zoo for fair comparison. As their adversarial instance perturbations cannot migrate to unseen query data, we compare the training data classification accuracy (Acc) in the cross-model setting. The results of them using the same model zoo in Tab. X show our method’s superiority and meta-learning’s effectiveness.
IV-D Adversarial Example Quality
To evaluate the image quality for generated adversarial examples, we compare the SSIM [79] with other attack methods for re-id. SSIM calculates structural similarity between synthetic and natural images and larger SSIM scores indicate better quality of synthetic images. The results of SSIM between AEs(=8) and benign images on DukeMTMC are show in Tab. XI, which shows that our MTGA can obtain AEs with comparable quality.
IV-E Visualization
We visualize the perturbations and adversarial examples that our MTGA generates. As Fig. 4 shows, the perturbations on adversarial examples are imperceptible. It’s hard for humans to detect the maliciously attacked adversarial examples generated by our MTGA. What’s more, the generated perturbations obtain the human shape of benign images, which indicates that our MTGA is able to understand the target that needs to be attacked and attempts to perform a full range of feature destruction for different person images, thus generating more generic adversarial attacks.
V Conclusion
In this paper, we propose a novel Meta Transferable Generative Attack method to facilitate the attacker generating highly transferable adversarial examples on black-box re-id models by learning from extensive simulated transfer-based meta attack tasks. The proposed Cross-model&dataset Attack Simulation method constructs the cross-model and cross-dataset attack tasks by selecting different model and data for meta-train and meta-test process. PRE strategy randomly erases the generated perturbation to suppress the model-specific feature corruption. NorMix module mimics diverse feature embeddings to boost the cross-test transferability. Comprehensive experiments show the superiority of our proposed MTGA over the state-of-the-art methods.